Copenhagen Covered (C). Heidi Hoi Jensen is again evaluating the arbitrage profit potential in the same market after another change in interest rates. (Remember that any time the difference in interest rates does not exactly equal the forward premium, it must be possible to make a CIA profit one way or another.)

Arbitrage funds available $5,000,000
Spot exchange rate (kr/$) 6.1720
3-month forward rate (kr/$) 6.1980
U.S. dollar 3-month interest rate 3.000%
Danish kroner 3-month interest rate 6.000%
Answer #1

A covered interest arbitrage profit can be made with these steps :

  • Borrow $5,000,000 at the US interest rate for 3 months. Convert into Kr at the spot exchange rate. Kr received = 5,000,000 * 6.1720 = Kr 30,860,000
  • Invest Kr 30,860,000 at the Kr interest rate for 3 months. Kr received after 3 months = Kr 30,860,000 * (1 + (6% * (3/12))) = Kr 31,322,900
  • Today, enter into a forward exchange contract to convert Kr 31,322,900 into $ after 3 months. $ received after 3 months =  Kr 31,322,900 / 6.1980 = $5,053,710.87
  • Repay the original $ loan. $ to repay = $5,000,000 * (1 + (3% * (3/12))) = $5,037,500
  • Arbitrage profit = $ received after 3 months - $ to repay after 3 months
  • Arbitrage profit = $5,053,710.87 - $5,037,500 = $16,210.87